A travel agency staffs its call center with seven agents. Calls arrive at a rate of forty per hour. The average time to service a call is ten minutes. If the agency pays its agents $30 per hour and the estimated cost to the customer is $20 per hour, what is the number of agents the travel agency should staff to deliver the lowest total cost to the system?
Answer #1

Here, Number of Servers = 7

Utilization= (40/60)/(7*1/10)=(40*10)/(60*7)=0.952

Average Number in the queue/Line = (0.952)^2/(1-0.952)=18.88

Wait in System=18.88+10=28.88

Average Number in the system=(28.88*40)/60=19.25

Total cost of the system per hour= (Average Number in the system*Cost to the customer per hour)+(Number of servers*Cost of servers per Hour)= (19.25*20)+(7*30)=$595 Hour